If you combine 240.0 mL of water at 25.00 ∘C and 130.0 mL of water at 95.00 ∘C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.
Answer #1

since density of water is 1g/mL
m1 = 240.0 g
T1 = 25.0 oC
m2 = 130.0 g
T2 = 95.0 oC

Let the final temperature be T oC

use:
heat lost by 2 = heat gained by 1
m2*c*(T2-T) = m1*c*(T-T1)
m2*(T2-T) = m1*(T-T1)
130.0*(95.0-T) = 240.0*(T-25.0)
T= 49.5946 oC
Answer: 49.59 oC